# Do LLMs Judge Distantly Supervised Named Entity Labels Well? Constructing the JudgeWEL Dataset

### v1

@@ Abstract
Alistair Plum, Laura Bernardy, Tharindu Ranasinghe
University of Luxembourg, Luxembourg
Lancaster University, UK
@@ 1.  Introduction
Recent advances in large language models (LLMs) have redefined the scope of natural language processing (NLP), shifting attention from classification and sequence-labelling tasks to increasingly complex generative applications. This transformation has enabled conversational systems that integrate reasoning, translation, and summarisation within a single multi-task architecture brown2020language; touvron2023llama; jiang2023mistral. Yet, the reach of these technologies remains uneven. For many less-resourced and under-represented languages, the benefits are limited: they continue to lack fundamental resources such as labelled corpora that support even basic supervised learning tasks, including sequence labelling and text classification joshi2020state; adelani2022masakhaner.
Building datasets for such languages remains particularly challenging. Obstacles typically include a scarcity of written material, limited linguistic expertise, and the absence of sustained funding, all of which reinforce a cycle of under-representation. When coupled with limited prior research, these constraints make it difficult to bootstrap the resources necessary for progress hettiarachchi-etal-2025-overview. In these circumstances, creative and data-efficient strategies for dataset construction are essential. This is the case for Luxembourgish, the national language of Luxembourg and one of three official languages alongside German and French. Despite its official status, Luxembourgish is the only national language of an EU member state not represented at the EU administrative level, thus remaining absent from standardised resources such as Europarl koehn2005europarl.
Luxembourgish is under-represented in NLP joshi2020state compared to its larger neighbours German and French, though it shares deep historical and structural ties with both. It evolved from the Moselle Franconian dialect Gilles2019, displaying strong affinities with German while integrating extensive French lexical and grammatical. Code-switching and borrowing are common, particularly in informal written communication, giving the language a dynamic character that further complicates computational treatment. Despite these challenges, the language’s roughly 400,000 speakers Gilles2019, ongoing standardisation efforts, and renewed digital interest make this an opportune moment to improve NLP for Luxembourgish.
@@
We hypothesise that certain LLMs can reliably distinguish high-quality annotated sentences from those containing missing or incorrect tags, thus enabling efficient filtering and scalable dataset construction. Our main contributions [^1] The dataset and code will be made publicly available after the review process. are as follows:
@@
    An open pipeline for automatically constructing annotated NER datasets using Wikipedia, Wikidata, and LLM-based judgements.
@@
    The judgeWEL dataset for Luxembourgish NER, consisting of 28,866 sentences featuring five entity types.
@@ 2.1.  Named Entity Recognition
NER is a fundamental Information Extraction and NLP task that identifies and classifies entities (e.g., people, locations, organizations, dates) via sequence labeling schemes such as BIO ramshaw-marcus-1995-text. It builds the base for applications like question answering, translation, summarization, and knowledge base construction.
Research on NER was started with MUC-6 sundheim1995overview. Early rule-based systems used linguistic heuristics mikheev1999named; weischedel1996progress and gazetteers rau1991extracting; farmakiotou2000rule; nadeau2006unsupervised, the latter forming the basis for knowledge-based methods exploiting ontologies mendes2011dbpedia; rizzo2012nerd. These approaches required no labelled data and training but lacked (cross-domain) scalability.
Recent work explores generative and LLM approaches. Zero-shot NER with LLMs as standalone remains challenging xie2023empiricalstudyzeroshotner; hu2024improvinglargelanguagemodels, but improvements are achieved through prompt engineering and few-shot learning with approaches like PromptNER ashok2023promptnerpromptingnamedentity, GL-NER 10.1007/978-3-031-72350-6_30, and ProML chen2022promptbasedmetriclearningfewshot. Hybrid frameworks like SuperICL xu2023smallmodelsvaluableplugins and LinkNER 10.1145/3589334.3645414 further integrate LLM reasoning with fine-tuned NER models to enhance accuracy.
@@
NER systems for high-resource (HR) languages such as English have achieved state-of-the-art performance due to the availability of large annotated datasets, including CoNLL’03 tjong2003_conll and OntoNotes weischedel2011ontonotes. However, reaching this performance is difficult in LR settings where labelled data is scarce. This lack of annotated data is often mitigated through data augmentation and knowledge transfer between HR and LR languages.
Cross-lingual knowledge transfer aims to transfer learned linguistic and semantic representations from a source language to a target language with limited labelled data. For example, mayhew-etal-2017-cheap used translation techniques based on the lexica and Wikipedia to map the knowledge from HR to LR languages. Similarly, ijcai2018p566 used an extended approach by enriching it with semantic mappings, enabling coverage of named entities unseen in the LR datasets.
tsai-roth-2016-cross developed a technique for cross-lingual Wikification, linking non-English Wikipedia entries to English Wikipedia, training multilingual embeddings based on it to facilitate cross-lingual transfer. The WikiAnn dataset pan-etal-2017-cross is a large multilingual benchmark covering 282 languages, also created via cross-lingual Wikipedia links. Likewise, Polyglot-NER leveraged distributed word representations encoding semantic and syntactic properties of words across languages, by integrating Wikipedia links and Freebase attributes to create a multilingual NER corpus in 40 languages doi:10.1137/1.9781611974010.66.
An intermediate approach between cross-lingual transfer and augmentation is MulDA liu-etal-2021-mulda. It addresses the absence of labelled data in a target language by translating limited NER data into multiple target languages using entity placeholders, then generating synthetic multilingual sequences through LMs.
For low-resource NER augmentation, PromptDA ijcai2022p590 introduced a label-conditioned word replacement and question-answering style prompting method using BERT. sym13050786 enhanced LR NER by fine-tuning XLM-RoBERTa using data from HR languages and an attention-based strategy to capture the most relevant semantic and syntactic informationMore recent augmentation methods employ LLMs for contextual and entity-level rewriting ye2024llmdadataaugmentationlarge, designed specifically for few-shot and low-resource settings.
Synthetic data generation via prompting LLMs with a small number of gold examples has also shown promising results kamath2025doessyntheticdatahelp. The application is further explored in studies examining the effectiveness of LLM-generated synthetic data for NER augmentation.kamath2025doessyntheticdatahelp
@@ 2.3.  Language Resources for Luxembourgish
Only recently has research on NLP for Luxembourgish gained momentum, with a few early exceptions marking the first steps in resource creation and linguistic analysis. adda-decker-etal-2008-developments presented initial tools and corpora for computational processing; snoeren-etal-2010-study investigated orthographic tendencies such as contextual n-deletion in transcribed speech; and lavergne-etal-2014-automatic compiled a manually annotated dataset of mixed-language sentences to test a language identification system. More recently, progress has been made with resources, tasks and models. sirajzade-etal-2020-sentiment and Gierschek2022 developed modern sentiment analysis pipelines, purschke2020attitudes proposed an automatic approach for orthographic correction, and plum2024 started the first treebank for Luxembourgish. philippy-etal-2024-forget introduced a zero-shot classification method based on a curated topic dictionary. ranasinghe-etal-2023-publish demonstrated a comment moderation task, lutgen2025 worked on orthographic normalization, and plum2025 introduced the generative benchmark set LuxGen.
@@
Various strategies have also been explored for Luxembourgish language models. These include transfer learning from German with LuxGPT bernardy2022, data augmentation via generated samples with LuxemBERT lothritz-etal-2022-luxembert2, and balanced training data extension via multilingual data for LuxT5 plum2025.
@@ 3.2.  Entity Linking
Linking entities to their corresponding Wikidata Vrandečić and Krötzsch (2014) page is the next step in the pipeline. Each article is split into sentences using the sentence-splitter Koehn and Schroeder (2005) package for Python. The splitter is not trained for Luxembourgish, but still performs best when set to German, compared to other options [^3] We manually tested a small set of Luxembourgish examples with other options for the language.. For each hyperlinked entity in a sentence, we match the article title via the Wikidata API to check its attributes. More specifically, we check the following attributes for each type of entity that we are targeting:
@@
If a hyperlinked entity matches any of the attributes corresponding to a targeted entity type, we annotate the entity tokens with that label. For the labelling, we adopt the BIO (Begin-Inside-Outside) encoding for entity boundary annotation, a format that has been widely used in NER ramshaw1995_chunking; tjong2003_conll. Any hyperlinked entity that does not match the targeted entity types, or for which we do not find a corresponding Wikidata entry, is discarded.
@@ 3.4.  Annotation Improvements
Since the Luxembourgish Wikipedia does not cover all entities present in our dataset due to its size, we perform an additional automatic annotation step. Using a LuxGPT model fine-tuned through prompt engineering for NER with a modified version of RTL-NER dataset bernardy2022, we re-evaluate all entities initially labelled with O-tags and, where necessary, reassign them with one of the predefined entity tags. Some manual post-processing is conducted to harmonize the tag set, including the unification of the GPE tag under LOC. This is because the RTL-NER distinguishes between these two tags (see Section 2.3, whereas we do not. These adjustments are implemented using simple regular expressions.
@@ 3.5.  LLM-as-a-Judge
GPT-5 represents the cutting edge of commercial models and serves as a reference benchmark for high-end proprietary performance (openai2025_introducing_gpt5). Its smaller counterpart, GPT-5-mini, provides a more lightweight and cost-efficient variant for comparison, enabling us to assess performance scalability within the same model family.
@@
Gemma-3-27B-IT is included because Google/DeepMind released it under a permissive licence, allowing transparent evaluation and reproducibility. Mistral-Medium-3.1 represents the new generation of mid-sized open models optimised for efficiency and multilingual generalisation, offering a valuable test case for open community-led development. Finally, GPT-OSS-120B acts as a large open-weight counterpart to the closed models, released by OpenAI under Apache 2.0 with near-parity on reasoning benchmarks, providing a direct comparison point for open vs. proprietary large-scale systems (openai2025_gptoss_model_card; openai2025_introducing_gpt_oss).
@@
LLaMA-3.3-8B-Instruct offers a smaller, instruction-optimised model within a widely adopted open-weight ecosystem (touvron2024_llama3). In addition, command-a-03-2025 represents a new generation of instruction-tuned commercial models that occupy the middle ground between fully open and closed systems, and is said to emphasise controlled outputs and multilingual understanding (cohere2025_command_a).
@@ 4.1.  LLM-as-a-Judge
Overall, GPT-5-mini and GPT-5 produce the exact same output, and clearly outperform all other systems, reaching a $\kappa$ value of 0.62, approaching the level of human inter-annotator agreement at 0.66. This indicates that high-end proprietary models can understand Luxembourgish text and capture the fine-grained cues that distinguish well-annotated from inconsistent examples. The fact that the smaller model produces the same output as the larger model also has significant cost implications, as the former cost roughly \$25 to annotate the entire pre-filtered set of 74k sentences, while the latter would have cost around \$180.
Among the open-weight systems, GPT-OSS-120B and Mistral-Medium-3.1 perform comparatively well, showing moderate and stable alignment with human judgements. The command-a-03-2025 model, representing the new generation of instruction-tuned commercial LLMs, achieves slightly lower agreement but remains consistent across samples, suggesting solid comprehension but a less decisive evaluative behaviour. By contrast, Gemma-3-27B-IT and LLaMA-3.3-8B-Instruct exhibit near-zero or negative correlation with human ratings, confirming that instruction tuning alone does not ensure evaluative reliability in smaller open-weight systems.
@@ 4.2.  Named Entity Recognition Task
For the NER task, we benchmark a set of multilingual and Luxembourgish-specific encoder models: mBERT devlin2018, LuxemBERT Lothritz et al. (2022), and XLM-RoBERTa conneau2019. Alongside this, we fine-tune and benchmark three generative LLMs: LuxGPT-NER, aya-expanse-8b 2024ayamodelinstructionfinetuned and Meta-Llama-3-8B-Instruct (touvron2024_llama3). We also tested GPT-5 in a zero- and few-shot setting, however, the model did not deliver consistent results (producing just O tags) and we do not discuss any further results [^4] Finding the optimal model for this task was not the objective of this research..
@@
To understand the potential performance increase achievable with judgeWEL, we carry out the evaluation in two settings. First, we train, validate and test on judgeWEL, constructed with the methodology laid out in this paper (Section 4.2.1). Second, we train and validate on judgeWEL, but test using the existing RTL-NER dataset for Luxembourgish (Section 4.2.2), which we previously described in Section 2.3.

### v2

@@ Abstract
Alistair Plum1, Laura Bernardy1, Tharindu Ranasinghe2
1University of Luxembourg, Luxembourg
2Lancaster University, UK
@@ 1.  Introduction
Recent advances in large language models (LLMs) have redefined the scope of natural language processing (NLP), shifting attention from classification and sequence-labelling tasks to increasingly complex generative applications. This transformation has enabled conversational systems that integrate reasoning, translation, and summarisation within a single multi-task architecture Brown et al. (2020); Touvron et al. (2023); Jiang et al. (2023). Yet, the reach of these technologies remains uneven. For many less-resourced and under-represented languages, the benefits are limited: they continue to lack fundamental resources such as labelled corpora that support even basic supervised learning tasks, including sequence labelling and text classification Joshi et al. (2020); Adelani et al. (2022).
Building datasets for such languages remains particularly challenging. Obstacles typically include a scarcity of written material, limited linguistic expertise, and the absence of sustained funding, all of which reinforce a cycle of under-representation. When coupled with limited prior research, these constraints make it difficult to bootstrap the resources necessary for progress Hettiarachchi et al. (2025). In these circumstances, creative and data-efficient strategies for dataset construction are essential. This is the case for Luxembourgish, the national language of Luxembourg and one of three official languages alongside German and French. Despite its official status, Luxembourgish is the only national language of an EU member state not represented at the EU administrative level, thus remaining absent from standardised resources such as Europarl Koehn (2005).
Luxembourgish is under-represented in NLP Joshi et al. (2020) compared to its larger neighbours German and French, though it shares deep historical and structural ties with both. It evolved from the Moselle Franconian dialect Gilles (2019), displaying strong affinities with German while integrating extensive French lexical and grammatical. Code-switching and borrowing are common, particularly in informal written communication, giving the language a dynamic character that further complicates computational treatment. Despite these challenges, the language’s roughly 400,000 speakers Gilles (2019), ongoing standardisation efforts, and renewed digital interest make this an opportune moment to improve NLP for Luxembourgish.
@@
We hypothesise that certain LLMs can reliably distinguish high-quality annotated sentences from those containing missing or incorrect tags, thus enabling efficient filtering and scalable dataset construction. Our main contributions are as follows:
@@
    An open pipeline for automatically constructing annotated NER datasets using Wikipedia, Wikidata, and LLM-based judgements. [^1] https://github.com/plumaj/judgeWEL
@@
    The judgeWEL dataset for Luxembourgish NER, consisting of 28,866 sentences featuring five entity types.1
@@ 2.1.  Named Entity Recognition
NER is a fundamental Information Extraction and NLP task that identifies and classifies entities (e.g., people, locations, organizations, dates) via sequence labelling schemes such as BIO Ramshaw and Marcus (1995a). It builds the base for applications like question answering, translation, summarization, and knowledge base construction.
Research on NER was started with MUC-6 Sundheim (1995). Early rule-based systems used linguistic heuristics Mikheev et al. (1999); Weischedel et al. (1996) and gazetteers Rau (1991); Farmakiotou et al. (2000); Nadeau et al. (2006), the latter forming the basis for knowledge-based methods exploiting ontologies Mendes et al. (2011); Rizzo and Troncy (2012). These approaches required no labelled data and training but lacked (cross-domain) scalability.
Recent work explores generative and LLM approaches. Zero-shot NER with LLMs as standalone remains challenging Xie et al. (2023); Hu et al. (2024); Premasiri et al. (2025), but improvements are achieved through prompt engineering and few-shot learning with approaches like PromptNER Ashok and Lipton (2023), GL-NER Zhu et al. (2024), and ProML Chen et al. (2023). Hybrid frameworks like SuperICL Xu et al. (2024) and LinkNER Zhang et al. (2024) further integrate LLM reasoning with fine-tuned NER models to enhance accuracy.
@@
NER systems for high-resource (HR) languages such as English have achieved state-of-the-art performance due to the availability of large annotated datasets, including CoNLL’03 Tjong Kim Sang and De Meulder (2003) and OntoNotes Weischedel et al. (2011). However, reaching this performance is difficult in LR settings where labelled data is scarce Ranasinghe et al. (2025). This lack of annotated data is often mitigated through data augmentation and knowledge transfer between HR and LR languages.
Cross-lingual knowledge transfer aims to transfer learned linguistic and semantic representations from a source language to a target language with limited labelled data. For example, Mayhew et al. (2017) used translation techniques based on the lexica and Wikipedia to map the knowledge from HR to LR languages. Similarly, Feng et al. (2018) used an extended approach by enriching it with semantic mappings, enabling coverage of named entities unseen in the LR datasets.
Tsai and Roth (2016) developed a technique for cross-lingual Wikification, linking non-English Wikipedia entries to English Wikipedia, training multilingual embeddings based on it to facilitate cross-lingual transfer. The WikiAnn dataset Pan et al. (2017) is a large multilingual benchmark covering 282 languages, also created via cross-lingual Wikipedia links. Likewise, Polyglot-NER leveraged distributed word representations encoding semantic and syntactic properties of words across languages, by integrating Wikipedia links and Freebase attributes to create a multilingual NER corpus in 40 languages Al-Rfou et al. (2015).
An intermediate approach between cross-lingual transfer and augmentation is MulDA Liu et al. (2021). It addresses the absence of labelled data in a target language by translating limited NER data into multiple target languages using entity placeholders, then generating synthetic multilingual sequences through LMs.
For low-resource NER augmentation, PromptDA Liu et al. (2022) introduced a label-conditioned word replacement and question-answering style prompting method using BERT. Chen et al. (2021) enhanced LR NER by fine-tuning XLM-RoBERTa using data from HR languages and an attention-based strategy to capture the most relevant semantic and syntactic information. More recent augmentation methods employ LLMs for contextual and entity-level rewriting Ye et al. (2024), designed specifically for few-shot and low-resource settings.
Synthetic data generation via prompting LLMs with a small number of gold examples has also shown promising results Kamath and Vajjala (2025). The application is further explored in studies examining the effectiveness of LLM-generated synthetic data for NER augmentation.Kamath and Vajjala (2025)
@@ 2.3.  Language Resources for Luxembourgish
Only recently has research on NLP for Luxembourgish gained momentum, with a few early exceptions marking the first steps in resource creation and linguistic analysis. Adda-Decker et al. (2008) presented initial tools and corpora for computational processing; Snoeren et al. (2010) investigated orthographic tendencies such as contextual n-deletion in transcribed speech; and Lavergne et al. (2014) compiled a manually annotated dataset of mixed-language sentences to test a language identification system. More recently, progress has been made with resources, tasks and models. Sirajzade et al. (2020) and Gierschek (2022) developed modern sentiment analysis pipelines, Purschke (2020) proposed an automatic approach for orthographic correction, and Plum et al. (2024) started the first treebank for Luxembourgish. Philippy et al. (2024) introduced a zero-shot classification method based on a curated topic dictionary. Ranasinghe et al. (2023) demonstrated a comment moderation task, Lutgen et al. (2025) worked on orthographic normalization, and Plum et al. (2025) introduced the generative benchmark set LuxGen.
@@
Various strategies have also been explored for Luxembourgish language models. These include transfer learning from German with LuxGPT Bernardy (2022), data augmentation via generated samples with LuxemBERT Lothritz et al. (2022), and balanced training data extension via multilingual data for LuxT5 Plum et al. (2025).
@@ 3.2.  Entity Linking
Linking entities to their corresponding Wikidata Vrandečić and Krötzsch (2014) page is the next step in the pipeline. Each article is split into sentences using the sentence-splitter Koehn and Schroeder (2005) package for Python. The splitter is not trained for Luxembourgish, but still performs best when set to German, compared to other options. [^3] We manually tested a small set of Luxembourgish examples with other options for the language. For each hyperlinked entity in a sentence, we match the article title via the Wikidata API to check its attributes. More specifically, we check the following attributes for each type of entity that we are targeting:
@@
If a hyperlinked entity matches any of the attributes corresponding to a targeted entity type, we annotate the entity tokens with that label. For the labelling, we adopt the BIO (Begin-Inside-Outside) encoding for entity boundary annotation, a format that has been widely used in NER Ramshaw and Marcus (1995b); Tjong Kim Sang and De Meulder (2003). Any hyperlinked entity that does not match the targeted entity types, or for which we do not find a corresponding Wikidata entry, is discarded.
@@ 3.4.  Annotation Improvements
Since the Luxembourgish Wikipedia does not cover all entities present in our dataset due to its size, we perform an additional automatic annotation step. Using a LuxGPT model fine-tuned through prompt engineering for NER with a modified version of RTL-NER dataset Bernardy (2022), we re-evaluate all entities initially labelled with O-tags and, where necessary, reassign them with one of the predefined entity tags. Some manual post-processing is conducted to harmonize the tag set, including the unification of the GPE tag under LOC. This is because the RTL-NER distinguishes between these two tags (see Section 2.3), whereas we do not. These adjustments are implemented using simple regular expressions.
@@ 3.5.  LLM-as-a-Judge
GPT-5 represents the cutting edge of commercial models and serves as a reference benchmark for high-end proprietary performance (OpenAI, 2025b). Its smaller counterpart, GPT-5-mini, provides a more lightweight and cost-efficient variant for comparison, enabling us to assess performance scalability within the same model family.
@@
Gemma-3-27B-IT is included because Google/DeepMind released it under a permissive licence, allowing transparent evaluation and reproducibility. Mistral-Medium-3.1 represents the new generation of mid-sized open models optimised for efficiency and multilingual generalisation, offering a valuable test case for open community-led development. Finally, GPT-OSS-120B acts as a large open-weight counterpart to the closed models, released by OpenAI under Apache 2.0 with near-parity on reasoning benchmarks, providing a direct comparison point for open vs. proprietary large-scale systems (OpenAI, 2025a, c).
@@
LLaMA-3.3-8B-Instruct offers a smaller, instruction-optimised model within a widely adopted open-weight ecosystem (Touvron et al., 2024). In addition, command-a-03-2025 represents a new generation of instruction-tuned commercial models that occupy the middle ground between fully open and closed systems, and is said to emphasise controlled outputs and multilingual understanding (Cohere for AI, 2025).
@@ 4.1.  LLM-as-a-Judge
Overall, GPT-5-mini and GPT-5 produce the exact same output, and clearly outperform all other systems, reaching a $\kappa$ value of 0.62, approaching the level of human inter-annotator agreement at 0.66. This indicates that high-end proprietary models can understand Luxembourgish text and capture the fine-grained cues that distinguish well-annotated from inconsistent examples. The fact that the smaller model produces the same output as the larger model also has significant cost implications, as the former cost roughly \$25 to annotate the entire pre-filtered set of 74k sentences, while the latter would have cost around \$180. Among the open-weight systems, GPT-OSS-120B and Mistral-Medium-3.1 perform comparatively well, showing moderate and stable alignment with human judgements. The command-a-03-2025 model achieves slightly lower agreement but remains consistent across samples, suggesting solid comprehension but a less decisive evaluative behaviour. By contrast, Gemma-3-27B-IT and LLaMA-3.3-8B-Instruct exhibit near-zero or negative correlation with human ratings, confirming that instruction tuning alone does not ensure evaluative reliability in smaller open-weight systems.
A finer-grained analysis of errors broken down by entity type reveals that these aggregate differences mask systematic per-category biases, evaluated on the 388 sentences for which both human annotators agreed. DATE-containing sentences are the most consistently handled across all models, with every system achieving its highest or near-highest F1 on this category, likely because temporal expressions provide reliable surface cues that LLM judges can exploit regardless of their general evaluative quality. PER entities show a wider spread: Mistral-Medium-3.1 and command-a-03-2025 both exceed F1 72, while GPT-OSS-120B drops to 39.2, driven by a strong conservative bias that causes it to discard 52 of the 71 gold-keep PER sentences, a pattern consistent with its overall tendency to over-reject. MISC entities prove the most problematic across the board: GPT-5 and GPT-OSS-120B achieve an F1 of zero, never retaining a MISC sentence that human annotators approved, while the remaining models reach at best 17, largely driven by false positives. This likely reflects the heterogeneous and language-specific nature of the MISC class in Luxembourgish NER. Finally, sentences containing no named entities expose the sharpest divide: GPT-5 and GPT-OSS-120B handle these near-perfectly (F1 99.2 and 97.4), while Gemma-3-27B-IT discards all 60 such sentences, yielding an F1 of zero. This reinforces the interpretation that its negative $\kappa$ reflects a systematic tendency to equate the absence of named entities with annotation failure, rather than random disagreement.
@@ 4.2.  Named Entity Recognition Task
For the NER task, we benchmark a set of multilingual and Luxembourgish-specific encoder models: mBERT Devlin et al. (2018), LuxemBERT Lothritz et al. (2022), and XLM-RoBERTa Conneau et al. (2019). Alongside this, we fine-tune and benchmark three generative LLMs: LuxGPT-NER, aya-expanse-8b Üstün et al. (2024) and Meta-Llama-3-8B-Instruct (Touvron et al., 2024). We also tested GPT-5 in a zero- and few-shot setting, however, the model did not deliver consistent results (producing just O tags) and we do not discuss any further results, as finding the optimal model for this task was not our objective.
@@
To understand the potential performance increase achievable with judgeWEL, we carry out the evaluation in two settings. First, we train, validate and test on judgeWEL (Section 4.2.1), constructed with the methodology laid out in this paper. Second, we train and validate on judgeWEL, but test using the existing RTL-NER dataset for Luxembourgish (Section 4.2.2), which we previously described in Section 2.3.